### v3

# Reconstruction the scalar-torsion gravity version from the frame of exact cosmological solutions

Abstract
We consider the cosmological models based on the scalar-torsion gravity implying non-minimal coupling between torsion and the scalar field as modification of teleparallel equivalent of general relativity. Based on generalized exact solutions of the equations of cosmological dynamics, the type of scalar-torsion gravity was reconstructed. Also, based on observational constraints on the values of the parameters of cosmological perturbations, the type of the coupling function was determined. It was noted that any models of cosmological inflation obtained on the basis of the proposed approach can be verified by observational constraints.
1 Introduction
Currently, a description of the inflationary stage of the evolution of the universe with including some inflaton fields for the case of general relativity and its various modifications (with and without inflaton fields) is necessary for constructing correct cosmological models [1, 2, 3]. The first works on inflationary cosmology [4, 5, 6, 7] taught their further development, and at the moment there are a large number of cosmological models describing both stage of accelerated expansion of the universe with different types of the scalar fields based on the Einstein gravity theory [1, 2], the teleparallel equivalent of the general relativity (TEGR) [8, 9] and their various extensions and modifications [3, 10, 11, 12, 13, 14, 15, 16, 17] are actively used.
Also, various mechanisms of the particle production are considered, including rapid oscillations of the inflaton near the minimum potential in accordance with the reheating process [18], instant preheating [19], curvaton reheating [20] and gravitational particle production [21], including the case of non-minimum coupling of torsion and material fields [22] as well.
Thus, at the moment, there is no specific theory of gravity or cosmological model that would be considered to unconditionally describe the evolution of the universe, and different cosmological models corresponding to observational data can be considered as actual ones.
The construction and analysis of cosmological models can be considered in two aspects, namely by the analysis of the possible physical effects induced by the different versions of the gravity theories [1, 2, 3], and, also, by the reconstruction of the model’s parameters from the various observational constraints including the results of measurements of anisotropy and polarization of CMB [23, 24, 25].
An important method for analyzing cosmological models is the use of the slow-roll approximation approach, which takes into account that the early universe at the inflationary stage expands close to exponential expansion law or close to the de Sitter stage [1, 2, 3]. However, exact cosmological solutions provide a clearer understanding of the specifics of inflationary processes or the parameters of cosmological models [26]. Methods for constructing exact cosmological solutions based on the different symmetries of cosmological dynamic equations, the nonlocal conservation laws, various form-invariant transformations and many other methods can be used to construct and analyze exact cosmological solutions (see, for example, in [2]). The background dynamics of the expansion of the universe and the evolution of cosmological perturbations can be considered on the basis of both exact and approximate cosmological solutions [1, 2, 3].
According to the inflationary paradigm, quantum fluctuations of the scalar field generate corresponding perturbations of the space-time metric, which are the source of large-scale structure and relict gravitational waves [1, 2, 3].
As a basis for analyzing the compliance of cosmological models with observational constraints on the values of the parameters of cosmological perturbations, the dependence of the tenor-to-scalar ratio on the spectral index of scalar perturbations $r=r(1-n_{S})$ is usually used to verify the cosmological models by observational constraints following from the results of measurements of anisotropy and polarization of CMB [27, 28].
At the moment, a large number of cosmological models corresponding to the observational constraints on the parameters of cosmological perturbations are considered [3, 11, 12, 13, 14, 15, 16]. For cosmological models based on general relativity (or TEGR), agreement with observational data will lead to an unambiguous determination of the model parameters including the potential of a scalar field [1, 2]. However, for the case of cosmological models based on modified theories of gravity, agreement with observational data can be obtained for ambiguously defined parameters [29, 30, 31, 32, 33, 34].
For example, in the case of cosmological models based on the scalar-tensor gravity theories [34, 35, 36, 37, 38, 39] in paper [34] it was shown that these models can meet observational constraints not due to the choice of model parameters, but due to a specific relationship between the parameters for arbitrary cosmological models.
Thus, within the framework of the approach proposed in [34], any cosmological model can be verified using observational data after finding the appropriate connection between the parameters of the modified theory of gravity and the parameters of the material source of gravity, which is considered to be scalar fields. As a criterion for finding such connections, the exact solutions of the cosmological dynamics equations and known types of scalar-tensor gravity and scalar field potentials were used [34].
In this paper we consider the application of this approach to the case of cosmological models based on the scalar-torsion gravity with non-minimal coupling of the scalar field and torsion, which generalize the case of the teleparallel equivalent of general relativity.
The description of the teleparallel gravity, its modifications including the non-minimal coupling of torsion and the scalar field, and their applications in cosmology were previously considered in many works (see, for example, in [40, 41, 42, 43]).
The properties of the cosmological models with scalar-torsion gravity expressed in a fairly general form $f(T,\phi)_{STG}=f(T,\phi)$ were considered earlier in [44, 45, 46] based on the slow-roll approximation. Also, specific cosmological models based on the exact solutions of dynamic equations were considered in [47].
This paper presents new generalized exact solutions of the cosmological dynamic equations for the certain type of scalar-torsion gravity $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$. Also, the relationship between the potential of the scalar field and the type of the coupling function between the scalar field and torsion was reconstructed. The exact solutions obtained are generalized both for arbitrary scalar field potentials (positive and negative due to the the signature reversal symmetry) [48, 49, 50, 51, 52, 53, 54], for arbitrary cosmological dynamics and for arbitrary evolution of scalar fields as well.
The first property of the cosmological models based on the proposed version of the scalar-torsion gravity is the possibility of their verification for an arbitrary model’s parameters for the case of a linear dependence of the tensor-to-scalar ratio on the spectral index of scalar perturbations $r\sim(1-n_{S})$, in contrast to the case of Einstein gravity (or TEGR) for which models with the dependence $r\sim(1-n_{S})$ are unverified by observational constraints [1, 2].
The second property of the proposed version of the scalar-torsion gravity is that relation $r\sim(1-n_{S})$ is preserved for an arbitrary type of cosmological dynamics, which implies an arbitrary choice of a scale factor $a=a(t)$ or Hubble parameter $H=H(t)$ for cosmological models verified by these observational constraints.
[…]
Thus, on the one hand, the proposed approach makes it possible to construct a large number of phenomenologically correct cosmological models, however, on the other hand, this property was obtained for the specific version of gravity theory, the discussion of which is devoted to this work.
The paper is organized as follows. In Sec. 2, the exact solutions of the cosmological dynamic equations for the scalar-torsion gravity of type $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$ are obtained, where $T$ is the torsion scalar, and function $G(\phi)$ defines the character of the non-minimal coupling between a scalar field and torsion. It is also shown that for the case of the quasi-de Sitter expansion of the universe, this type of the scalar-torsion gravity is close to the case of the teleparallel equivalent of general relativity. In Sec. 3, expressions for the parameters of cosmological perturbations are obtained in explicit form for the inflation based on the scalar-torsion gravity under consideration. Sec. 4 discusses a new classification of cosmological models according to the order of expansion of the dependence of the tensor-to-scalar ratio on the spectral index of scalar perturbations $r=r(1-n_{S})$. Relationships between the Hubble parameter $H$, the non-minimal coupling function $G$ and the scalar field potential $V$ are obtained for cosmological models based on the proposed version of scalar-torsion gravity, which are verified by observational constraints on the values of the parameters cosmological perturbations in the first order of this expansion $r\sim(1-n_{S})$. It is also shown that, taking into account these relationships, cosmological models with arbitrary dynamics of the expansion of the universe (subject to the slow-roll conditions at the inflationary stage) based on the scalar-torsion gravity $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$ will correspond to observational constraints on the values of the parameters of cosmological perturbations. In Sec. 5, as an example of the proposed approach to the construction the cosmological models based on scalar-torsion gravity $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$, the case of the exponential scalar field potential is considered. It is shown that for this case it is possible to reconstruct exact expressions for all other parameters of cosmological models. Also, using the examples of well-known cases of exponential power-law inflation [58, 59, 60, 61, 62] and Quasi-Steady State cosmological model [63, 64, 65, 66], it is shown that the proposed gravity theory allows both to describe the first and second accelerated expansion of the universe and to propose a solution to the Hubble tension problem with taking into account the observational constraints on the parameters of cosmological perturbations. Sec. 6 discusses the general properties of cosmological models based on the proposed version of scalar-torsion gravity in comparison with models based on the other modified gravity theories. Our results are summarized in Conclusion.
2 Cosmological models in generalized scalar-torsion gravity
The generalized action for cosmological models based on the scalar-torsion gravity with non-minimal coupling of a scalar field $\phi$ and torsion $T$ in the system of units $c=8\pi G=M^{-2}_{p}=1$ can be written as [44, 45, 46]
$S=\int d^{4}x\,e\,F(T,X,\phi)=\int d^{4}x\,e\,\left[f(T,\phi)+\omega(\phi)X\right],$ (1)
where $f=f(T,\phi)$ is an arbitrary differentiable function of a scalar field $\phi$ and torsion scalar $T$, also kinetic energy of a scalar field $X=-\partial_{\mu}{\phi}\partial^{\mu}{\phi}/2$.
The action (1) includes non-minimally coupled scalar-torsion gravity models with $f(T,\phi)$ and the kinetic function $\omega=\omega(\phi)$.
In choosing the cosmological background, we assume the diagonal tetrad field
$e^{A}_{~{}\mu}={\rm diag}(1,a,a,a),$ (2)
which is a proper tetrad naturally associated with the vanishing spin connections $\omega^{A}_{~{}B\mu}=0$, and which leads to the spatially flat Friedmann-Robertson-Walker (FRW) metric [44, 45, 46]
$ds^{2}=-dt^{2}+a^{2}\,\delta_{ij}dx^{i}dx^{j}\,,$ (3)
where $a=a(t)$ is the scale factor and $t$ is the cosmic time.
The background equations for this model can be written as [44, 45, 46]
$f(T,\phi)-\frac{1}{2}\omega(\phi)\dot{\phi}^{2}-2Tf_{,T}=0,$ (4)
$f(T,\phi)+\frac{1}{2}\omega(\phi)\dot{\phi}^{2}-2Tf_{,T}-4\frac{d}{dt}\left(Hf_{,T}\right)=0,$ (5)
$\omega(\phi)\ddot{\phi}+3\omega(\phi)H\dot{\phi}+\frac{1}{2}\frac{d\omega(\phi)}{d\phi}\dot{\phi}^{2}-f_{,\phi}=0,$ (6)
where $H\equiv\dot{a}/a$ is the Hubble rate, and a dot represents derivative with respect to $t$. Partial derivatives is denoted as $f_{,\phi}=\frac{\partial f}{\partial\phi},~{}f_{,T}=\frac{\partial f}{\partial T}$.
It should be noted, that the torsion scalar is $T=6H^{2}$ for the case of the spatially flat Friedmann-Robertson-Walker metric [44, 45, 46].
From here, using the relation $H=\sqrt{\frac{T}{6}}$ and taking into account the condition $H>0$, we can rewrite the dynamic equations (4)–(6) as
$f(T,\phi)-\frac{1}{2}\omega(\phi)\dot{\phi}^{2}-2Tf_{,T}=0,$ (7)
$\frac{1}{2}\omega(\phi)\dot{\phi}^{2}-2\frac{d}{dt}\left(\sqrt{\frac{T}{6}}f_{,T}\right)=0,$ (8)
$\omega(\phi)\ddot{\phi}+3\omega(\phi)\sqrt{\frac{T}{6}}\dot{\phi}+\frac{1}{2}\frac{d\omega(\phi)}{d\phi}\dot{\phi}^{2}-f_{,\phi}=0.$ (9)
Such representation of dynamic equations allow us select the class of models, admitting exact solutions of the equations (7)–(9) for arbitrary Hubble parameter $H=H(t)$.
The exact solutions of the system of equations (7)–(9) are represented by the special form of the function $f(T,\phi)$, connected with the potential $V(\phi)$ and kinetic term $\frac{1}{2}\omega(\phi)\dot{\phi}^{2}$ and can be written as
$f(T,\phi)=\alpha_{1}G(\phi)\sqrt{T}+\alpha_{2}V(\phi)=f(T,\phi)_{STG}+\alpha_{2}V(\phi),$ (10)
$\omega(\phi)=\frac{\alpha^{2}_{1}}{3\alpha_{2}}\frac{G_{,\phi}^{2}}{V(\phi)},$ (11)
$\dot{\phi}=\frac{\sqrt{6}\alpha_{2}}{\alpha_{1}}\frac{V(\phi)}{G_{,\phi}},$ (12)
where $G(\phi)$ is arbitrary differentiable function of a scalar field.
The function
$f(T,\phi)_{STG}=\alpha_{1}G(\phi)\sqrt{T},$ (13)
fixed the version of the scalar-torsion gravity and $\alpha_{1}$, $\alpha_{2}$ are the normalization constants, which allow one to consider the function (10) with different signs.
Using expression $\ddot{\phi}=\frac{d\dot{\phi}}{d\phi}\dot{\phi}$, one can verify by direct substitution, that solutions (10)–(12) satisfy all three dynamic equations (7)–(9).
Let us note, that by using $G_{,\phi}=\dot{G}/\dot{\phi}$ expression (12) can be rewritten as follow
$\dot{G}=\frac{\sqrt{6}\alpha_{2}}{\alpha_{1}}\,V(\phi).$ (14)
Thus, the expressions (10)–(14) define the frame of exact solutions of cosmological dynamic equations (7)–(9) for the inflationary models based on scalar-torsion gravity for arbitrary Hubble parameter $H=H(t)$.
As it was noted in [45], for the function
$f(T,\phi)=-\frac{1}{2}T-V(\phi),$ (15)
one has the case of TEGR, and dynamic equations (7)–(9) for the case $\omega=1$ are reduced to well known ones
$3H^{2}=\frac{1}{2}\dot{\phi}^{2}+V(\phi),$ (16)
$\dot{\phi}^{2}=-2\dot{H},$ (17)
$\ddot{\phi}+3H\dot{\phi}+V_{,\phi}=0,$ (18)
which are the same ones for inflationary models based on the Einstein gravity [1, 2].
Since at the stage of cosmological inflation the rate of expansion of the universe is close to exponential regime (or close to the de Sitter stage) $T\simeq T_{0}=6H^{2}_{0}=const$, expanding $\sqrt{T}$ into a series around $T=T_{0}$ we get
$\sqrt{T}=\frac{1}{2}\sqrt{T_{0}}+\frac{1}{2\sqrt{T_{0}}}T+\mathcal{O}\left[(T-T_{0})^{2}\right].$ (19)
Thus, for quasi de Sitter inflationary stage from (10) and (19) we obtain
$f(T,\phi)\simeq\alpha_{1}\tilde{G}(\phi)T+\alpha_{2}\tilde{V}(\phi),$ (20)
where $\tilde{G}(\phi)=\frac{G(\phi)}{2\sqrt{T_{0}}}$ and $\tilde{V}(\phi)=V(\phi)+T_{0}\tilde{G}(\phi)$.
Therefore, at the inflationary stage, exact expression for the function (10) can be reduced to the case of TERG (15) for $\tilde{G}(\phi)=-1/(2\alpha_{1})=const$, $\alpha_{2}=-1$ with fairly high accuracy.
Nevertheless, to describe the gravitational interaction we will consider the function (13) in general form as exact solutions consequence of cosmological dynamics equations. Besides, the form of a coupling function $G=G(\phi)$ in expression (13) is not determined.
Now, based on the nature of early universe accelerated expansion (close to exponential one) and observational constraints on the values of cosmological perturbations parameters, we reconstruct the connection between the coupling function $G=G(\phi)$ and the scalar field potential $V=V(\phi)$. For this purpose, let us consider cosmological perturbations in inflationary models based on the scalar-torsion gravity and generalized exact solutions (10)–(14) of cosmological dynamic equations (7)–(9).
3 Cosmological perturbations
In accordance with the theory of cosmological perturbations, quantum fluctuations of the scalar field induce the corresponding perturbations of the space-time metric during the inflationary stage. In the linear order of cosmological perturbation theory, the observed anisotropy and polarization of cosmic microwave background radiation (CMB) [1, 2] are explained by the influence of two types of perturbations, namely, scalar and tensor ones.
Observational constraints on the parameters of cosmological perturbations due to the modern observations of the anisotropy and polarisation of CMB are [27, 28]
$P_{S}=2.1\times 10^{-9},$ (21)
$n_{S}=0.9663\pm 0.0041,$ (22)
$r<0.032.$ (23)
Cosmological perturbations for the inflationary models based on the scalar-torsion gravity and action (1) were considered on the basis of the Arnowitt-Deser-Misner (ADM) decomposition of the tetrad field in order to define the perturbed metrics in [45]. Further, based on a standard analysis of the equations for the evolution of cosmological perturbations, expressions for the parameters of cosmological perturbations were obtained (see, for details, in [45]).
The expressions for the parameters of cosmological perturbations at the crossing of the Hubble radius ($k=aH$) for the cosmological models based on action (1) can be written as follows [45]
$\mathcal{P}_{S}=\frac{H^{2}}{8\pi^{2}Q_{S}}\left[1+2\eta_{\mathcal{R}}\ln\left(\frac{k}{aH}\right)\right]_{k=aH}=\frac{H^{2}}{8\pi^{2}Q_{S}},$ (24)
$\mathcal{P}_{T}=\frac{H^{2}}{2\pi^{2}Q_{T}},$ (25)
$n_{S}-1=-2\epsilon-\eta+2\eta_{\mathcal{R}},$ (26)
$n_{T}=-2\epsilon-\delta_{f_{,T}},$ (27)
$r=\frac{\mathcal{P}_{T}}{\mathcal{P}_{S}}=16\delta_{\omega X},$ (28)
where $\mathcal{P}_{S}$ and $\mathcal{P}_{T}$ are the power spectra of scalar and tensor perturbations, $n_{S}$ and $n_{T}$ are the spectral index of the scalar and tensor perturbations and $r$ is the tensor-to-scalar ratio, also, $n_{T}<0$ corresponds to the red tilt tensor spectrum, $n_{T}>0$ corresponds to the blue tilt tensor spectrum and $n_{T}=0$ corresponds to the flat tensor spectrum.
Also, the parameters $Q_{S}$ and $Q_{T}$ are [45]
$Q_{S}=\frac{\omega X}{H^{2}},~{}~{}~{}Q_{T}=-\frac{1}{2}f_{,T},$ (29)
and the slow-roll parameters in expressions (24)–(28) are defined as follows
$\epsilon=-\frac{\dot{H}}{H^{2}},~{}~{}~{}~{}\delta_{\omega X}=-\frac{\omega X}{2H^{2}f_{,T}},~{}~{}~{}~{}\delta_{f_{,T}}=\frac{\dot{f}_{,T}}{Hf_{,T}},$ (30)
$\delta_{f\dot{H}}=\frac{f_{,TT}\dot{T}}{Hf_{,T}},~{}~{}~{}\delta_{fX}=\frac{f_{,T\phi}\dot{\phi}}{Hf_{,T}},~{}~{}~{}\eta_{\mathcal{R}}=\delta_{f_{,T}}\left[1+\frac{\delta_{f_{,T}}}{\delta_{f\dot{H}}}\left(1+\frac{\delta_{fX}}{\delta_{\omega X}}\right)\right],$ (31)
$\eta=\frac{\dot{Q}_{S}}{HQ_{S}}=3\epsilon+\frac{1}{H}\left[\frac{d}{dt}\ln(\omega XH)\right]=2\epsilon+\frac{1}{H}\left(\frac{\dot{\omega}}{\omega}+\frac{\dot{X}}{X}\right),$ (32)
where all these parameters much less than unity ($\epsilon\ll 1$) due to the quasi de Sitter nature of the inflationary stage of the evolution of the early universe.
On the basis of expressions (11)–(12) and (14) for the proposed exact solutions we obtain following expression
$\omega X=\frac{\alpha_{1}}{\sqrt{6}}\dot{G}=\alpha_{2}V.$ (33)
Thus, from (21), (24), (29) and (33) we obtain following condition
$\mathcal{P}_{S}=\left(\frac{H^{2}}{8\pi^{2}Q_{S}}\right)_{k=aH}=\left(\frac{H^{4}}{8\pi^{2}\alpha_{2}V}\right)_{k=aH}=2.1\times 10^{-9}.$ (34)
Thus, form expression (34) we have the following conditions: $\alpha_{2}>0$ for the positive potentials $V>0$ and $\alpha_{2}<0$ for the negative potentials $V<0$ of a scalar field.
The examples of different cosmological models with positive and negative potentials of a scalar field one can find in [48, 49, 50, 51, 52, 53, 54].
Since condition (34) can always be satisfied by choosing the parameters of the inflationary model, as a criterion for verifying cosmological models we will consider the dependence of the tensor-scalar ratio on the spectral index of scalar perturbations $r=r(1-n_{S})$.
Taking into account (33) and expression $T=6H^{2}$, for the inflationary models described by solutions (10)–(12) from expressions (30)–(32) we obtain
$\delta_{\omega X}=\frac{\sqrt{6}\,\omega X}{HG}=\frac{\alpha_{1}\dot{G}}{HG},~{}~{}~{}~{}\delta_{f_{,T}}=\epsilon+\delta_{fX},~{}~{}~{}~{}\delta_{f\dot{H}}=\epsilon,$ (35)
$\delta_{fX}=\frac{\sqrt{6}\alpha_{2}}{\alpha_{1}}\,\frac{V}{HG}=\frac{\dot{G}}{HG}=\frac{\delta_{\omega X}}{\alpha_{1}},~{}~{}~{}~{}~{}~{}~{}\eta=2\epsilon+\frac{\ddot{G}}{H\dot{G}},$ (36)
$\eta_{\mathcal{R}}=\delta_{f_{,T}}\left[1+\left(1+\frac{\delta_{fX}}{\epsilon}\right)\left(1+\frac{1}{\alpha_{1}}\right)\right]=$
$=\left(\epsilon+\frac{\dot{G}}{HG}\right)\left[1+\left(1-\frac{H\dot{G}}{G\dot{H}}\right)\left(1+\frac{1}{\alpha_{1}}\right)\right].$ (37)
Thus, we can write the spectral indices of the scalar and tensor perturbations and the tensor-to-scalar ratio as
$n_{S}-1=4\frac{\dot{H}}{H^{2}}-\frac{\ddot{G}}{H\dot{G}}+2\left(-\frac{\dot{H}}{H^{2}}+\frac{\dot{G}}{HG}\right)\left[1+\left(1-\frac{H\dot{G}}{G\dot{H}}\right)\left(1+\frac{1}{\alpha_{1}}\right)\right],$ (38)
$r=16\alpha_{1}\frac{\dot{G}}{HG},$ (39)
$n_{T}=3\frac{\dot{H}}{H^{2}}-\frac{\dot{G}}{HG}=-3\epsilon-\frac{r}{16\alpha_{1}}.$ (40)
As one can see from expression (40), for $\alpha_{1}>-\frac{r}{48\epsilon}$ one has red tilt tensor spectrum, for $\alpha_{1}<-\frac{r}{48\epsilon}$ one has blue tilt tensor spectrum, and for $\alpha_{1}=-\frac{r}{48\epsilon}$ one has flat tensor spectrum.
4 Reconstruction of non-minimal coupling function from linear dependence $r\sim(1-n_{S})$
Now, we will determine the connection between the coupling function $G$ and the Hubble parameter $H$, for dependence $r=r(1-n_{S})$ in the main (linear) order and consider quasi de Sitter solutions for the model under consideration.
Since the value of the spectral index of scalar perturbations is $n_{S}\simeq 0.97$ and $1-n_{S}\simeq 0.03\ll 1$, we can write the dependence $r=r(1-n_{S})$ as follows
$r=\sum^{\infty}_{k=0}\beta_{k}(1-n_{S})^{k}=\beta_{0}+\beta_{1}(1-n_{S})+\beta_{2}(1-n_{S})^{2}+...,$ (41)
where $(1-n_{S})$ is the small parameter of expansion and $\beta_{k}$ are the constant coefficients.
Since, the zeroth order term in this expansion $r(0)=\beta_{0}=0$ from condition $r(n_{S}=1)=0$ corresponding to the flat Harrison-Zel’dovich spectrum [67, 68], we can rewrite expression (41) in the following form
$r=\sum^{\infty}_{k=1}\beta_{k}(1-n_{S})^{k}=\beta_{1}(1-n_{S})+\beta_{2}(1-n_{S})^{2}+...$ (42)
Since the first term in expression (42) makes the main contribution to the value of the tensor-scalar ratio, we can consider the dependence $r=r(1-n_{S})$ with a reasonable degree of accuracy in the first order on the basis of expressions (38)–(39).
Thus, we can consider the new specific classification of cosmological models (regardless of gravity type or cosmological model’s parameters) according to non-zero orders of expansion of dependence $r=r(1-n_{S})$ as follows: first-order models for $\beta_{1}\neq 0$, second-order models for $\beta_{1}=0$ and $\beta_{2}\neq 0$ etc.
For the purpose to obtain the connection between parameters of the cosmological models under consideration for the first-order models in explicit form with arbitrary Hubble parameter (i.e. for arbitrary cosmological dynamics), we consider the special case $\alpha_{1}=-1$, and from expressions (38)–(39) we obtain
$n_{S}-1=2\frac{\dot{H}}{H^{2}}+2\frac{\dot{G}}{HG}-\frac{\ddot{G}}{H\dot{G}},$ (43)
$r=-16\frac{\dot{G}}{HG}.$ (44)
The linear dependence between tensor-to-scalar ratio and spectral index of scalar perturbations we define as follows
$r\simeq\frac{8}{m}(1-n_{S}),$ (45)
where $\beta_{1}=8/m$ and $m>0$ is the positive constant.
Also, we note that dependence (45) is explicit one and it doesn’t depend on the background dynamic parameters, i.e. this dependence is valid for any parameters of the cosmological model, including the times corresponding to the e-folds number $N\approx 60$ at the end of the inflationary stage [1, 2].
Also, on the basis of the observational constraints $n_{S}\simeq 0.97$ and $r<0.032$ from (45) for verifiable inflationary models one has following constraint $m>7.5$.
Figure 1: The dependence $r=r(n_{S})$ for different values of the parameter $m=7,8,12,24$ with constraints on the tensor-to-scalar ratio $r$ due to the Planck TT,TE,EE+lowE+lensing+BAO+BICEP2/Keck Array observations [27, 28].
Figure 1 shows dependence (45) for various values of $m$. It can be seen that under condition $m>7.5$, the values of the spectral index of scalar perturbations $n_{S}$ and the tensor-scalar ratio $r$ fall into the internal region of allowed values, which corresponds to the possibility of verifying the inflationary model according to the observational constraints.
From equations (43)–(44) and (45) we can reconstruct corresponding connection between Hubble parameter $H$ and coupling function $G$ for this case, namely
$H(t)=const\times\dot{G}^{1/2}G^{m-1}.$ (46)
From equations (14) and (46) for such a models we finally obtain
$H(t)=const\times V^{1/2}G^{m-1},$ (47)
where the Hubble parameter $H>0$ and has the real values for any type of potential (positive or negative) due to the fact that the constant in expression (47) is arbitrary one.
For the pure de Sitter regime of accelerated expansion of the early universe $H=const$ from (47) we obtain
$G(\phi)=const\times\left[V(\phi)\right]^{-\frac{1}{2(m-1)}},$ (48)
for any type of the scalar field evolution $\phi=\phi(t)$, where $G_{0}$ is a some arbitrary constant.
Therefore, taking into account quasi de Sitter regime of accelerated expansion of the early universe $H\simeq const$, we can represent the exact expression for the coupling function as follows
$G(\phi)=const\times\left[V(\phi)\right]^{-\frac{1+k}{2(m-1)}},$ (49)
where parameter $k$ defines the deviations from the pure de Sitter regime.
We also note that expression (49) is one of the possible ways to parameterize deviations from the de Sitter stage in expression for the coupling function, and we use it as a fairly convenient one to analyze exact solutions of the cosmological dynamics equations for the certain inflationary models.
Thus, the gravitational term in action (1) for exact solutions (10)–(14) in quasi de Sitter regime $H\simeq const$ can be defined as
$f(T,\phi)_{STG}=-G(\phi)\sqrt{T}=cosnt\times\sqrt{T}\times\left[V(\phi)\right]^{-\frac{1+k}{2(m-1)}},$ (50)
i.e. the type of the non-minimal coupling of a scalar field $\phi$ with torsion scalar $T$ depends on the type of its potential.
After substitution of expression (49) into (47) we obtain
$H(t)=const\times\left[V(\phi)\right]^{-k/2}.$ (51)
Also, for $\alpha_{1}=-1$, from expression $\alpha_{1}=-\frac{r}{48\epsilon}$ we obtain the following condition on the value of the slow-roll parameter $\epsilon=r/48$ corresponding to the flat tensor spectrum, for $\epsilon>r/48$ one has red tilted tensor spectrum, and for $480<\epsilon<r/$ 48 one has blue tilted tensor spectrum. Thus, these inflationary models can predict different types of tensor perturbation spectrum depending on how close the cosmological dynamics of the early universe are to the purely exponential (de Sitter) expansion regime.
5 Inflation with exponential potential based on the scalar-torsion gravity $f(T,\phi)_{STG}=-G(\phi)\sqrt{T}$
As the example of cosmological model implying generalized scalar field potential (in the sense of positive and negative values), the exit from inflation and the second accelerated expansion of the universe we consider the model with exponential potential, which was considered earlier in the context of inflationary models based on GR, namely
$V(\phi)=V_{0}\exp(-2\beta\phi),$ (52)
where $\beta$ is the positive constant, and $V_{0}$ can take positive and negative values [54], that can be considered as consequence of the signature reversal symmetry [50].
From equations (11), (49) and (51) for potential (52) we obtain expressions for the Hubble parameter, coupling function, and kinetic function in explicit form
$H(t)=\lambda\exp\left(\beta k\phi\right),$ (53)
$G(\phi)=G_{0}\exp\left(\frac{\beta(k+1)}{m-1}\phi\right),$ (54)
$\omega(\phi)=\omega_{0}\exp\left(\frac{2\beta(m+k)}{m-1}\phi\right),$ (55)
where $\lambda$, $G_{0}$ and $\omega_{0}$ are a some constants.
Also, from expression (53) and following representation the Hubble parameter
$H(t)=\lambda+f(t),$ (56)
we obtain explicit relation between the type of the scalar field evolution and cosmological dynamics
$\phi(t)=\frac{1}{\beta k}\ln\left(1+\frac{f(t)}{\lambda}\right),$ (57)
where $f=f(t)$ is a some function of cosmic time.
Thus, in this case, we can construct exact solutions with arbitrary cosmological dynamics for first-order models based on the scalar-torsion gravity under consideration.
5.1 Exponential-power law inflation
As the example of the function corresponding to inflationary stage, the graceful exit from inflation and the second accelerated expansion of the universe at the large times we consider the case $f(t)=s/t$, where $s$ is a some constant.
The Hubble parameter and the scale factor in this case are
$H(t)=\lambda+\frac{s}{t},$ (58)
$a(t)\sim t^{s}\exp(\lambda t),$ (59)
where $s$ is a some positive constant, and $\dot{H}=-s/t^{2}<0$, i.e the Hubble parameter is the decreasing function of cosmic time.
Now, we define the e-fold number for this model as follows
$N(t)=\ln\left(\frac{a}{a_{0}}\right)=\lambda t+s\ln t,$ (60)
where $a_{0}$ is the scale factor at the beginning of inflation corresponding to $N=0$.
From expression (60) we can obtain the inverse dependence
$t(N)=\exp\left[\frac{N}{s}-W\left(\frac{\lambda}{s}e^{N/s}\right)\right],$ (61)
where $W$ is denote the Lambert function.
After substituting the dependence (61) into (57) and (52) we obtain the scalar field and the potential as the functions of the e-folds number
$\phi(N)=\frac{1}{\beta k}\ln\left\{1+\frac{s}{\lambda}\exp\left[W\left(\frac{\lambda}{s}e^{N/s}\right)-\frac{N}{s}\right]\right\},$ (62)
$V(N)=V_{0}\left\{1+\frac{s}{\lambda}\exp\left[W\left(\frac{\lambda}{s}e^{N/s}\right)-\frac{N}{s}\right]\right\}^{-2/k}.$ (63)
As one can see, the scalar field and potential at the beginning of inflation $\phi_{i}=\phi(N=0)$, $V_{i}=V(N=0)$ and at the end of inflation $\phi_{e}=\phi(N\approx 60)$, $V_{e}=V(N\approx 60)$ can take on different values due to the fact that they are determined by the five constant parameters of this model $\beta$, $k$, $s$, $\lambda$ and $V_{0}$.
We also note that the energy scale of inflation, which is defined as $E_{f}=|V|^{1/4}$ [69, 70] can be expressed as $E_{f}=|V_{0}e^{-2\beta\phi}|^{1/4}\sim 10^{-2}$ in chosen system of units or $E_{f}=|V_{0}e^{-2\beta\phi}|^{1/4}\sim 10^{-2}M_{P}\sim 10^{16}$ GeV [1, 69, 70] gives the additional relation between constant model’s parameters.
A detailed analysis of cosmological dynamics, within the framework of which it was shown that scale factor (59) implies an accelerated inflationary expansion of the universe, exit from inflation and second inflation (with a much lower rate of accelerated expansion of the universe than during the first inflation) at large times, was presented in [58, 59, 60, 61, 62].
We only note the general property of models with function $f=f(t)$ decreasing in time (as for the case of the function $f(t)=s/t$).
If condition $f(t\rightarrow\infty)=0$ is met, then from expressions (52)–(57) at the large times we have $\phi=0$, $H=\lambda$, $G=G_{0}$, $\omega=\omega_{0}$ and $V=V_{0}$.
Thus, at large times one has $f(t\gg t_{inf})\ll\lambda$ and $H\simeq\lambda$, and from (20) we obtain the model closed to the teleparallel equivalent of general relativity
$f(T,\phi)\simeq-\frac{G_{0}}{2\sqrt{T_{0}}}T+\alpha_{2}\left(V_{0}+\frac{1}{2}G_{0}\sqrt{T_{0}}\right)=$
$-\frac{1}{2}T+\alpha_{2}\left(V_{0}+\frac{1}{2}G^{2}_{0}\right)=-\frac{1}{2}T-\Lambda,$ (64)
with the cosmological constant $\Lambda=-\alpha_{2}\left(V_{0}+\frac{1}{2}G^{2}_{0}\right)$ for $G_{0}=\sqrt{T_{0}}$.
We also note that taking into account the observational value of the cosmological constant, we can consider the negative potential $V_{0}<0$ (and $\alpha_{2}<0$ due to condition (34), for the partial case $\alpha_{2}=-1$) to obtain $\Lambda=-\alpha_{2}\left(V_{0}+\frac{1}{2}G^{2}_{0}\right)=(2.846\pm 0.076)\times 10^{-122}$ [27].
Obviously, to build a correct model of the post-inflationary universe, it is necessary to take into account additional material fields, in particular, baryonic matter, radiation and Dark Matter. However, since we have a cosmological constant at the large times only, for the Cold Dark Matter this model at present era can be reduces to one close to the $\Lambda$CDM model based on TEGR.
5.2 Cyclic universe
We also note that based on the proposed approach, it is possible to obtain exact solutions for an arbitrary types of cosmological dynamics, for example, for a cyclic universe or Quasi-Steady State cosmological model, with the following Hubble parameter [63, 64, 65, 66]
$H(t)=\lambda-\frac{\mu\alpha\sin(\mu t)}{1+\alpha\cos(\mu t)},~{}~{}~{}f(t)=-\frac{\mu\alpha\sin(\mu t)}{1+\alpha\cos(\mu t)},$ (65)
and scale factor
$a(t)\sim\exp(\lambda t)\left(1+\alpha\cos(\mu t)\right),$ (66)
where constant $\mu>0$ and $10\leq\alpha<$ 1 in order to obtain the positive scale factor $a(t)>0$ for any cosmic time.
In this case, the Hubble parameter might increase at late times to account the well-known tension between PLANCK and local observations [55, 56, 57].
Corresponding scalar field evolution (57) for the Hubble parameter (65) is
$\phi(t)=\frac{1}{\beta k}\ln\left[1-\frac{\mu\alpha\sin(\mu t)}{\lambda(1+\alpha\cos(\mu t))}\right].$ (67)
In this case, the potential (52), coupling function (54) and kinetic function (55) can be written as follows
$V(t)=V_{0}\left[1-\frac{\mu\alpha\sin(\mu t)}{\lambda(1+\alpha\cos(\mu t))}\right]^{-2/k},$ (68)
$G(t)=G_{0}\left[1-\frac{\mu\alpha\sin(\mu t)}{\lambda(1+\alpha\cos(\mu t))}\right]^{\frac{k+1}{k(m-1)}},$ (69)
$\omega(t)=\omega_{0}\left[1-\frac{\mu\alpha\sin(\mu t)}{\lambda(1+\alpha\cos(\mu t))}\right]^{\frac{2(m+k)}{k(m-1)}},$ (70)
i.e. they are the periodic functions of cosmic time, and model is reduced to the case of cosmological constant $\Lambda$ and TEGR at the times closed to $t_{\Lambda}=\pi n/\mu$, where $n$ are a some positive integer.
Thus, the Hubble tension in this case can be explained by the deviation of the observation time in present era from this value $t_{obs}=\frac{\pi}{\mu}(n+\Delta)$.
The solution to the Hubble tension problem based on the Quasi-Steady State cosmological model based on background dynamics without analyzing the model’s compliance with observational constraints on the anisotropy and polarization of CMB was discussed earlier in [65, 66]. In the case of the proposed model, these constraints are satisfied due to relation (45). Other possible solutions to the Hubble tension problem can be found, for example, in [55, 56, 57].
Obviously, the cosmological models with other scalar field potentials and different types of cosmological dynamics will also satisfy observational constraints on the values of the parameters of cosmological perturbations due to dependence (45) for the proposed gravity theory.
6 The properties of the proposed models
As one can see, different cosmological dynamics does not affect the linear relation $r\sim(1-n_{S})$ between tensor-to scalar ratio and spectral tilt of the scalar perturbations (45), i.e. the cosmological models under consideration with different dynamics of accelerated expansion of the early universe can be considered as verifiable ones by the observational constraints on the parameters of cosmological perturbations.
Also, an obvious generalization of the proposed approach is the analysis of cosmological models based on the other scalar field potentials which differ from exponential one (52) based on expressions (11), (49) and (51) or for representations of the coupling function other than (49).
Thus, the proposed scalar-torsion gravity $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$ allows one to consider inflationary models with an arbitrary potential of the scalar field and arbitrary cosmological dynamics (satisfying the correct change of stages of expansion of the universe) as in principle verifiable by observational constraints already at the first (linear) order of expansion of dependence $r=r(1-n_{S})$.
Let us note the results presented in [34], in which the verification of inflationary models based on scalar-tensor gravity also did not depend on the type of scalar field potential or the specific type of scalar-tensor gravity, however, in this case, the quadratic dependence $r\sim(1-n_{S})^{2}$ corresponding to the second-order models within the framework of our classification was obtained.
However, we note that verification of inflationary models with different dependencies $r=r(1-n_{S})$ with specific model’s parameters can be achieved as for the case of GR [48] and for the case of the other modified gravity theories as well (see, for example, in [29, 30, 31, 32, 33]).
For these reasons, the proposed approach is one of the many possible ways to construct the cosmological models. Nevertheless, the main advantage of the proposed approach is the ability to obtain the verified cosmological models for an arbitrary model’s parameters (including arbitrary dynamics of the expansion of the universe) for the exactly defined type of the scalar-torsion gravity. This property of the proposed theory of gravity that makes it attractive for further research and applications to actual problems of cosmology and astrophysics.
7 Conclusion
In this work, cosmological models based on scalar-torsion gravity with non-minimal coupling between the scalar field and torsion, which was obtained as consequence of the generalized exact solutions of cosmological dynamic equations for an arbitrary Hubble parameter $H=H(t)$ and an arbitrary evolution of a scalar field $\phi=\phi(t)$.
Based on these generalized exact solutions, it was concluded that the gravitational interaction within the framework of scalar-torsion gravity in such models should be determined by a function $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$, which corresponds to some corrections to the usual teleparallel equivalent of general relativity $f(T,\phi)_{TEGR}\sim T$. It was also noted that at the inflationary quasi de Sitter stage of the evolution of the universe, these corrections are quite small.
Further, based on the previously proposed approach to analyzing the evolution of cosmological perturbations for the cosmological models based on the scalar-torsion gravity for general case $f(T,\phi)_{STG}=f(T,\phi)$, we determined the parameters of cosmological perturbations for the case of the considered version of scalar-tensor gravity $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$ on the basis of exact background solutions of cosmological dynamic equations.
A classification of inflation models according to the order of expansion of the dependence tensor-to-scalar ratio from spectral index of the scalar perturbations $r=r(1-n_{S})$ was also proposed, and on the basis of this classification the method for constructing verified by observational constraints inflationary models based on scalar-torsion gravity $f(T,\phi)_{STG}=-G(\phi)\sqrt{T}$ was considered, implying a linear dependence $r\sim(1-n_{S})$ for arbitrary model’s parameters with specific relation (46) between the Hubble parameter and the coupling function. We note that in this case, the type of scalar field potential $V=V(\phi)$ or other background parameters do not affect the possibility of verifying the proposed cosmological models.
It was also noted that proposed inflationary models can predict different types of tensor perturbation spectrum (red, blue of flat) depending on the value o the tensor-to-scalar ratio and how close the cosmological dynamics of the early universe are to the purely exponential (de Sitter) expansion regime.
Also, due to the correspondence of arbitrary models of cosmological inflation based on the generalized exact solutions to observational constraints on the parameters of cosmological perturbations, the following relationship between the non-minimal coupling function and the potential of the scalar field $G(\phi)\sim\left[V(\phi)\right]^{-\frac{1+k}{2(m-1)}}$ for quasi de Sitter dynamics accelerated expansion of the universe with the Hubble parameter $H\simeq const$ was obtained. Thus, the type of non-minimal coupling between the scalar field and torsion depends on the type of the scalar field potential.
As an example, we consider cosmological models based on exponential potential. It was shown that by choosing the Hubble parameter $H=H(t)$ it is possible to construct models with an exit from inflation and the second accelerated expansion of the universe in the present era. Also, as a possible solution to the Hubble tension problem, a model of a cyclic universe was considered on the basis of the proposed scalar-torsion gravity theory.
In conclusion, we note that proposed type of the scalar-torsion gravity $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$ reconstructed from exact solutions of cosmological dynamic equations and implying the wide class of verified cosmological models with arbitrary parameters is of interest for the further research of the deviations in the spectrum of relict gravitational waves and in the evolution of compact astrophysical objects from teleparallel equivalent of general relativity or from the other modified gravity theories.